OpenStudy (anonymous):

12. CD on a line from D= -5 1/4 and b=2/3 it says find the lengh. im in geometry 11th grade

OpenStudy (anonymous):

please specify your question more specifically

OpenStudy (anonymous):

from what in gather from your questio, perhaps you want to find C first, then you will be able to get the distance or lentgh of CD

OpenStudy (anonymous):

it on a number line and on he left side its negative 5 1/4 and on the right side its positive 2/3 so d=-5 1/4 and b =2/3 i cant draw it my comp. wont let me

OpenStudy (anonymous):

okay, can you take the differnce of those two line, th should give you the midpoint or c

OpenStudy (anonymous):